# Diversity or Precision? A Deep Dive into Next Token Prediction

### v1

@@ 1 Introduction
Recent advancements have demonstrated that reinforcement learning (RL) (bai2022rlhf; guo2025deepseekr1) can significantly enhance the reasoning capabilities of large language models (LLMs) (deepmind2025gemini; guo2025deepseekr1; anthropic2025claude; kimi2025kimik2). By utilizing verifiable rewards, such as passing unit tests or deriving correct mathematical solutions, LLMs evolve from merely mimicking human data to actively searching for optimal reasoning paths (guo2025deepseekr1). On-policy training paradigms have proven effective in unlocking the potential of pre-trained LLMs, prompting researchers to investigate how token output distributions influence RL. Recent studies (wang2025beyond82rule; zhu2025negrl; cui2025entropyrl; gandhi2025cogbeh) indicate that uncertainty in chain-of-thought reasoning is concentrated within a small subset of high-entropy forking tokens that govern pivotal decisions, while the majority of tokens exhibit low entropy. This observation underscores the critical impact of the pre-trained model’s output distribution on subsequent RL outcomes.
@@
Concurrently, researchers have explored next-token and next-segment reasoning objectives to derive self-supervised signals from massive unlabeled pre-training corpora (zelikman2024quietstar; dong2025rpt; li2025rlpt; xing2025pretrainzero). Applying RL to the pre-training corpus suggests a theoretical bridge connecting pre-training and RL. Specifically, next-token prediction can be reformulated as a reasoning task optimized via RL algorithms, where the model receives verifiable rewards for accurately predicting the subsequent token according to a given context. Notably, if the intermediate reasoning process is omitted, resulting in the direct generation of the answer, this procedure becomes analogous to standard pre-training. From the perspective of policy optimization, next-token prediction serves a foundational role by defining the initial policy distribution for subsequent RL. This distribution establishes the model’s behavioral trajectory and implicitly constrains its exploration space, thereby determining which reasoning paths the model prioritizes during RL.
@@
Motivated by this connection, we revisit the cross-entropy loss for next token prediction. Although traditionally viewed as a supervised metric, cross-entropy can be interpreted as a specific instance of policy gradient optimization within a single-step episode (wu2025dft; ming2025otr). This interpretation suggests that next-token prediction inherently permits an on-policy perspective, even though standard teacher forcing utilizes off-policy samples drawn directly from the training corpus distribution. From an entropy perspective, cross-entropy implicitly assigns maximal reward to the single ground-truth token while uniformly suppressing all negative tokens. Building on this insight, we aim to establish a unified pre-training objective that subsumes cross-entropy as a special case, enabling a systematic study of how reward configurations during pre-training influence subsequent RL dynamics.
@@ 2.1 Next Token Prediction
Building upon Equation 3, we treat the generation of a single token as a complete episode (ming2025otr). The objective for a fixed state $s_{t}$ simplifies to:
@@ 3.1 Training Settings
The training pipeline proceeds in three stages: pre-training, mid-training, and RLVR. Adhering to the Qwen3 (yang2025qwen3), we develop LLMs using both dense and MoE architectures. Specifically, we develop a series of LLMs, which include 1B and 4B dense models, as well as 5B-A0.3B and 10B-A0.5B MoE models. Moreover, we conduct the complete training pipeline on the 4B and 10B-A0.5B models, while the 1B and 5B-A0.3B models undergo the pre-training stage only. More training details are provided in Appendix A and Appendix B.
@@ 3.2 Evaluation Settings
  • General Knowledge: MMLU (hendrycks2020mmlu)(4-shot, CoT), MMLU-Pro (wang2024mmlupro)(5-shot, CoT), TriviaQA (joshi2017triviaqa)(5-shot), and NaturalQuestions (kwiatkowski2019naturalquestions)(5-shot).
  • Commonsense Reasoning: Hellaswag (zellers2019hellaswag)(0-shot), SIQA (sap2019socialiqa)(0-shot), PIQA (bisk2020piqa)(0-shot), WinoGrande (sakaguchi2021winogrande)(0-shot), OpenBookQA (mihaylov2018openbookqa)(5-shot), and CommonsenseQA (talmor2018commonsenseqa)(5-shot)
  • Logic Reasoning: ARC-Easy (clark2018arc)(0-shot), ARC-Challenge (clark2018arc)(0-shot), and BBH (suzgun2022bbh)(3-shot, CoT)
  • Mathematics: GSM8K (cobbe2021gsm8k)(4-shot, CoT), MATH-500 (lightman2023math500)(4-shot, CoT), Minerva (lewkowycz2022minerva)(4-shot, CoT), and OlympiadBench (he2024olympiadbench)(0-shot).
  • Coding: HumanEval+ (liu2023evalplus)(0-shot) and MBPP+ (liu2023evalplus)(3-shot).
Specifically, general knowledge and commonsense reasoning evaluate the model’s knowledge-base capabilities, whereas logical reasoning, mathematics, and coding probe its reasoning-base capabilities. Moreover, we employ the $\text{Pass@}k$ metric to evaluate the model’s upper-bound capability for tasks requiring mathematical reasoning and code generation. Pass@k measures the probability that at least one correct solution is present within $k$ independent attempts. We utilize the unbiased estimator of $\text{Pass@}k$ (chen2021passk), which is defined as:
@@
Evaluation of RL Models. For RL models evaluation, we employ various mathematics benchmarks, including AMC23 (maa2025amc), AIME (maa2025aime), MATH-500 (lightman2023math500), Minerva (lewkowycz2022minerva), and OlympiadBench (he2024olympiadbench). We sample 128 responses per problem and report Avg@128, Cons@128, and Pass@64 metrics. Specifically, $\text{Avg@}128$ represents the average accuracy across all 128 samples, while $\text{Cons@}128$ refers to the majority voting accuracy. Similarly, we configure the maximum output length to 16K for RL models.
@@
Figure 2: Changes of PPL and entropy during pre-training across 5B-A0.3B and 10B-A0.5B MoE models, developed based on different configurations.
@@ 3.3 Pre-Training
Our analysis of the proposed generalized training objective reveals that it effectively regulates the trade-off between diversity and precision by strategically varying reward configurations. As illustrated in Figure 1 and Figure 2, perplexity (PPL) consistently converges to comparable low values across both dense (1B, 4B) and MoE (5B-A0.3B, 10B-A0.5B) architectures. This demonstrates that, within a specific range, modifying the reward function modulates training dynamics without compromising final predictive accuracy. The parameter $\beta$ serves as a potent global entropy regulator. Specifically, setting $\beta<0$ significantly reduces entropy, resulting in a more peaked and confident token distribution by amplifying rewards for ground turth tokens. Conversely, $\beta>0$ maintains higher entropy and a flatter distribution, thereby promoting diversity in the generated output. Meanwhile, the parameters $\hat{\lambda}$ and $\tilde{\lambda}$ facilitate local entropy fine-tuning. These parameters shape the token distribution by either rewarding ($\hat{\lambda}=0,\tilde{\lambda}=0.1,k=100$) or penalizing ($\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$) negative tokens, enabling granular control over the training process.
@@
Figure 4: Changes of performance during mid-training across 4B dense and 10B-A0.5B MoE models, developed based on different configurations.
@@ 3.4 Mid-Training
3.5 Reinforcement Learning
Finally, we investigate the performance dynamics during the RL training stage across various actor models, as illustrated in Figure 5 and Figure 6. Pre-trained models derived from different reward configurations exhibit distinct output distributions, leading to significant variations in subsequent RL and end-to-end reasoning performance. Regarding the global entropy regulator $\beta$, we observe a consistent and robust trend across both the 4B dense and 10B-A0.5B MoE models. Specifically, the global low entropy setting ($\beta=-0.25$) yields superior performance trajectories. This configuration consistently outperforms the global high entropy setting across all evaluated metrics, including Avg@128, Cons@128, and Pass@64. Furthermore, the configuration $\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$ demonstrates a significant advantage, consistently achieving the highest performance and notably surpassing the baseline on the 10B-A0.5B MoE model. For the 4B dense model, maintaining local high entropy exhibits a superior scaling trend compared to the baseline. In conclusion, strategies that promote precision, either globally via generously rewarding positive tokens or locally by aggressively penalizing tail negative tokens, enables the model to converge to higher-quality solutions, potentially providing a better exploration space for RL.
@@ 3.5 Reinforcement Learning
Figure 7: Changes of entropy and response length during RL training across various actor models, developed based on 4B dense and 10B-A0.5B MoE architectures under different configurations.
@@ 3.6 $\text{Pass@}k$ Analysis of Base Models
Figure 8: $\text{Pass@}k$ curve of base models on mathematics reasoning and code generation tasks, developed based on 4B dense and 10B-A0.5B MoE models under different configurations.
@@ 4.1 Weighted Cross-Entropy Loss
The standard cross-entropy objective can be generalized within a policy-gradient framework, where it is equivalent to optimizing a sparse reward defined as $r_{\text{CE}}(s_{t},a_{t})=\mathbbm{1}(a_{t}=x_{t})\pi_{\theta}(a_{t}\mid s_{t})^{-1}$. Existing modifications to this objective include smooth loss (label smoothing), which encourages diversity by allocating a uniform probability mass to all positive tokens, and focal loss (lin2018focalloss), which down-weights easy examples via $w_{t}=(1-\pi_{\theta}(x_{t}\mid s_{t}))^{\gamma}$. Our proposed generalized training objective can also formulate these established variations. In this paper, we specifically explore two different reward configurations within this framework. Firstly, we introduce a modified positive reward, which is equivalent to applying a state-dependent weight $w_{t}=\pi_{\theta}(x_{t}\mid s_{t})^{1-(1-\pi_{\theta}(x_{t}\mid s_{t}))^{\beta}}$ to the standard cross-entropy. In addition, we incorporate TopK-based negative shaping, which explicitly controls local entropy by assigning non-zero rewards to selected actions with $a_{t}\neq x_{t}$.
@@ 4.2 Next Token Reasoning
Treating each token emission as a distinct episode ensures that the reward depends only on the immediate state-action pair $(s_{t}, a_{t})$, thereby preserving unbiased credit assignment.  The framework is naturally compatible with architectures that perform iterative internal computation prior to token emission, including latent-reasoning models (zelikman2024quietstar) and loop transformers (dehghani2019universaltransformers; zhu2025ouro). Although each episode terminates at token emission, the state $s_{t}$ may encode the outcome of internal refinement cycles. Our reward design can serve as an uncertainty-aware learning signal that can be combined with adaptive computation policies to allocate additional internal processing steps in uncertain contexts. By explicitly shaping positive and negative token-level rewards within a single-step policy-gradient framework, we provide a general and controllable mechanism that natively supports reasoning-oriented architectures through principled reward design.
@@ A.1 Implementation Details
For both the pre-training and mid-training phases, we employ the AdamW (loshchilov2017adamw) optimizer, implementing a weight decay of 0.1 and applying gradient clipping at 1.0. Throughout these stages, we utilize a warmup-stable-decay learning rate schedule with a global batch size of 16M. During the stable pre-training stage, which encompasses 500B tokens, the learning rate warms up over 2000 steps before stabilizing at $3\times 10^{-4}$. Subsequently, we perform mid-training on an additional 100B tokens, gradually decaying the learning rate from $3\times 10^{-4}$ to $3\times 10^{-5}$. We set the maximum sequence length to 4096 during pre-training and extend it to 16384 for the mid-training stage. To support long-context modeling during mid-training, we increase the base frequency of RoPE (su2024roformer) from $1e^{4}$ to $1e^{6}$.
@@ A.2 Model Architecture
Building upon the Qwen3 (yang2025qwen3) architectures, we perform our experiments utilizing both dense and MoE architectures. Notably, we adopt an auxiliary loss free approach (liu2024deepseekv3) for the training of the MoE models. Detailed architecture settings are provided in Table 1, where $E$ denotes the total number of experts and $E_{a}$ denotes the number of active experts.
@@ A.3 Experiment Results
We report comprehensive evaluation results to demonstrate performance progression throughout the training process. Tables 2 to 9 present the pre-training results across various models and different training tokens. Similarly, Tables 10 to 13 summarize the performance metrics for the mid-training stage.
@@ B.1 Implementation Details
For RLVR on mathematical reasoning tasks, we employ the on-policy GRPO algorithm (shao2024deepseekmath) without KL regularization. Following yu2025dapo, we incorporate clip-higher and dynamic sampling strategies to stabilize training. The process is conducted in two stages: an initial 700 steps with a sequence length of 8K, followed by continued training at a sequence length of 16K. We maintain a batch size of 128 and a constant learning rate of $1\times 10^{-6}$ for two stages. During training, we sample 16 outputs per prompt at a temperature of 1.0.

### v2

@@ 1 Introduction
Recent advancements have demonstrated that reinforcement learning (RL) (Bai et al., 2022; Guo et al., 2025) can significantly enhance the reasoning capabilities of large language models (LLMs) (Google DeepMind, 2025; Guo et al., 2025; Anthropic, 2025; Kimi et al., 2025). By utilizing verifiable rewards, such as passing unit tests or deriving correct mathematical solutions, LLMs evolve from merely mimicking human data to actively searching for optimal reasoning paths (Guo et al., 2025). On-policy training paradigms have proven effective in unlocking the potential of pre-trained LLMs, prompting researchers to investigate how token output distributions influence RL. Recent studies (Wang et al., 2025; Zhu et al., 2025b; Cui et al., 2025; Gandhi et al., 2025) indicate that uncertainty in chain-of-thought reasoning is concentrated within a small subset of high-entropy forking tokens that govern pivotal decisions, while the majority of tokens exhibit low entropy. This observation underscores the critical impact of the pre-trained model’s output distribution on subsequent RL outcomes.
@@
Concurrently, researchers have explored next-token and next-segment reasoning objectives to derive self-supervised signals from massive unlabeled pre-training corpora (Zelikman et al., 2024; Dong et al., 2025; Li et al., 2025; Xing et al., 2025). Applying RL to the pre-training corpus suggests a theoretical bridge connecting pre-training and RL. Specifically, next-token prediction can be reformulated as a reasoning task optimized via RL algorithms, where the model receives verifiable rewards for accurately predicting the subsequent token according to a given context. Notably, if the intermediate reasoning process is omitted, resulting in the direct generation of the answer, this procedure becomes analogous to standard pre-training. From the perspective of policy optimization, next-token prediction serves a foundational role by defining the initial policy distribution for subsequent RL. This distribution establishes the model’s behavioral trajectory and implicitly constrains its exploration space, thereby determining which reasoning paths the model prioritizes during RL.
@@
Motivated by this connection, we revisit the cross-entropy loss for next token prediction. Although traditionally viewed as a supervised metric, cross-entropy can be interpreted as a specific instance of policy gradient optimization within a single-step episode (Wu et al., 2025; Ming et al., 2025). This interpretation suggests that next-token prediction inherently permits an on-policy perspective, even though standard teacher forcing utilizes off-policy samples drawn directly from the training corpus distribution. From an entropy perspective, cross-entropy implicitly assigns maximal reward to the single ground-truth token while uniformly suppressing all negative tokens. Building on this insight, we aim to establish a unified pre-training objective that subsumes cross-entropy as a special case, enabling a systematic study of how reward configurations during pre-training influence subsequent RL dynamics.
@@ 2.1 Next Token Prediction
Building upon Equation 3, we treat the generation of a single token as a complete episode (Ming et al., 2025). The objective for a fixed state $s_{t}$ simplifies to:
@@ 3.1 Training Settings
The training pipeline proceeds in three stages: pre-training, mid-training, and RLVR. Adhering to the Qwen3 (Yang et al., 2025), we develop LLMs using both dense and MoE architectures. Specifically, we develop a series of LLMs, which include 1B and 4B dense models, as well as 5B-A0.3B, 10B-A0.5B and 20B-A1B MoE models. Moreover, we conduct the complete training pipeline on the 4B, 10B-A0.5B and 20B-A1B models, while the 1B and 5B-A0.3B models undergo the pre-training stage only. More training details are provided in Appendix A and Appendix B.
@@ 3.2 Evaluation Settings
  • General Knowledge: MMLU (Hendrycks et al., 2020)(4-shot, CoT), MMLU-Pro (Wang et al., 2024)(5-shot, CoT), TriviaQA (Joshi et al., 2017)(5-shot), and NaturalQuestions (Kwiatkowski et al., 2019)(5-shot).
  • Commonsense Reasoning: Hellaswag (Zellers et al., 2019)(0-shot), SIQA (Sap et al., 2019)(0-shot), PIQA (Bisk et al., 2020)(0-shot), WinoGrande (Sakaguchi et al., 2021)(0-shot), OpenBookQA (Mihaylov et al., 2018)(5-shot), and CommonsenseQA (Talmor et al., 2018)(5-shot)
  • Logic Reasoning: ARC-Easy (Clark et al., 2018)(0-shot), ARC-Challenge (Clark et al., 2018)(0-shot), and BBH (Suzgun et al., 2022)(3-shot, CoT)
  • Mathematics: GSM8K (Cobbe et al., 2021)(4-shot, CoT), MATH-500 (Lightman et al., 2023)(4-shot, CoT), Minerva (Lewkowycz et al., 2022)(4-shot, CoT), and OlympiadBench (He et al., 2024)(0-shot).
  • Coding: HumanEval+ (Liu et al., 2023)(0-shot) and MBPP+ (Liu et al., 2023)(3-shot).
Specifically, general knowledge and commonsense reasoning evaluate the model’s knowledge-base capabilities, whereas logical reasoning, mathematics, and coding probe its reasoning-base capabilities. Moreover, we employ the $\text{Pass@}k$ metric to evaluate the model’s upper-bound capability for tasks requiring mathematical reasoning and code generation. Pass@k measures the probability that at least one correct solution is present within $k$ independent attempts. We utilize the unbiased estimator of $\text{Pass@}k$ (Chen, 2021), which is defined as:
@@
Evaluation of RL Models. For RL models evaluation, we employ various mathematics benchmarks, including AMC23 (MAA, b), AIME (MAA, a), MATH-500 (Lightman et al., 2023), Minerva (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). We sample 128 responses per problem and report Avg@128, Cons@128, and Pass@64 metrics. Specifically, $\text{Avg@}128$ represents the average accuracy across all 128 samples, while $\text{Cons@}128$ refers to the majority voting accuracy. Similarly, we configure the maximum output length to 16K for RL models.
@@
Figure 2: Changes of PPL and entropy during pre-training across 5B-A0.3B, 10B-A0.5B and 20B-A1B MoE models, developed based on different configurations.
@@ 3.3 Pre-Training
Our analysis of the proposed generalized training objective reveals that it effectively regulates the trade-off between diversity and precision by strategically varying reward configurations. As illustrated in Figure 1 and Figure 2, perplexity (PPL) consistently converges to comparable low values across both dense (1B, 4B) and MoE (5B-A0.3B, 10B-A0.5B, 20B-A1B) architectures. This demonstrates that, within a specific range, modifying the reward function modulates training dynamics without compromising final predictive accuracy. The parameter $\beta$ serves as a potent global entropy regulator. Specifically, setting $\beta<0$ significantly reduces entropy, resulting in a more peaked and confident token distribution by amplifying rewards for ground turth tokens. Conversely, $\beta>0$ maintains higher entropy and a flatter distribution, thereby promoting diversity in the generated output. Meanwhile, the parameters $\hat{\lambda}$ and $\tilde{\lambda}$ facilitate local entropy fine-tuning. These parameters shape the token distribution by either rewarding ($\hat{\lambda}=0,\tilde{\lambda}=0.1,k=100$) or penalizing ($\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$) negative tokens, enabling granular control over the training process.
@@
Figure 4: Changes of performance during mid-training across 4B dense, 10B-A0.5B MoE and 20B-A1B MoE models, developed based on different configurations.
@@ 3.4 Mid-Training
Figure 7: Changes of performance during RL training across various actor models, developed based on a 20B-A1B MoE architecture under different configurations.
3.5 Reinforcement Learning
Finally, we investigate the performance dynamics during the RL training stage across various actor models, as illustrated in Figures 5 to 7. Pre-trained models derived from different reward configurations exhibit distinct output distributions, leading to significant variations in subsequent RL and end-to-end reasoning performance. Regarding the global entropy regulator $\beta$, we observe a consistent and robust trend across both the 4B dense and 10B-A0.5B MoE models. Specifically, the global low entropy setting ($\beta=-0.25$) yields superior performance trajectories. This configuration consistently outperforms the global high entropy setting across all evaluated metrics, including Avg@128, Cons@128, and Pass@64. Furthermore, the configuration $\hat{\lambda}=-0.1,\tilde{\lambda}=0,k=100$ demonstrates a significant advantage, consistently achieving the highest performance and notably surpassing the baseline on the 10B-A0.5B MoE model. For the 4B dense model, maintaining local high entropy exhibits a superior scaling trend compared to the baseline. In conclusion, strategies that promote precision, either globally via generously rewarding positive tokens or locally by aggressively penalizing tail negative tokens, enables the model to converge to higher-quality solutions, potentially providing a better exploration space for RL.
Figure 8: Changes of entropy and response length during RL training across various actor models, developed based on 4B dense, 10B-A0.5B MoE, and 20B-A1B MoE architectures under different configurations.
@@ 3.6 $\text{Pass@}k$ Analysis of Base Models
Figure 9: $\text{Pass@}k$ curve of base models on mathematics reasoning and code generation tasks, developed based on 4B dense, 10B-A0.5B MoE and 20B-A1B MoE models under different configurations.
@@ 4.1 Weighted Cross-Entropy Loss
The standard cross-entropy objective can be generalized within a policy-gradient framework, where it is equivalent to optimizing a sparse reward defined as $r_{\text{CE}}(s_{t},a_{t})=\mathbbm{1}(a_{t}=x_{t})\pi_{\theta}(a_{t}\mid s_{t})^{-1}$. Existing modifications to this objective include smooth loss (label smoothing), which encourages diversity by allocating a uniform probability mass to all positive tokens, and focal loss (Lin et al., 2018), which down-weights easy examples via $w_{t}=(1-\pi_{\theta}(x_{t}\mid s_{t}))^{\gamma}$. Our proposed generalized training objective can also formulate these established variations. In this paper, we specifically explore two different reward configurations within this framework. Firstly, we introduce a modified positive reward, which is equivalent to applying a state-dependent weight $w_{t}=\pi_{\theta}(x_{t}\mid s_{t})^{1-(1-\pi_{\theta}(x_{t}\mid s_{t}))^{\beta}}$ to the standard cross-entropy. In addition, we incorporate TopK-based negative shaping, which explicitly controls local entropy by assigning non-zero rewards to selected actions with $a_{t}\neq x_{t}$.
@@ 4.2 Next Token Reasoning
Treating each token emission as a distinct episode ensures that the reward depends only on the immediate state-action pair $(s_{t}, a_{t})$, thereby preserving unbiased credit assignment.  The framework is naturally compatible with architectures that perform iterative internal computation prior to token emission, including latent-reasoning models (Zelikman et al., 2024) and loop transformers (Dehghani et al., 2019; Zhu et al., 2025a). Although each episode terminates at token emission, the state $s_{t}$ may encode the outcome of internal refinement cycles. Our reward design can serve as an uncertainty-aware learning signal that can be combined with adaptive computation policies to allocate additional internal processing steps in uncertain contexts. By explicitly shaping positive and negative token-level rewards within a single-step policy-gradient framework, we provide a general and controllable mechanism that natively supports reasoning-oriented architectures through principled reward design.
@@ A.1 Implementation Details
For both the pre-training and mid-training phases, we employ the AdamW (Loshchilov & Hutter, 2017) optimizer, implementing a weight decay of 0.1 and applying gradient clipping at 1.0. Throughout these stages, we utilize a warmup-stable-decay learning rate schedule with a global batch size of 16M. During the stable pre-training stage, which encompasses 500B tokens, the learning rate warms up over 2000 steps before stabilizing at $3\times 10^{-4}$. Subsequently, we perform mid-training on an additional 100B tokens, gradually decaying the learning rate from $3\times 10^{-4}$ to $3\times 10^{-5}$. We set the maximum sequence length to 4096 during pre-training and extend it to 16384 for the mid-training stage. To support long-context modeling during mid-training, we increase the base frequency of RoPE (Su et al., 2024) from $1e^{4}$ to $1e^{6}$.
@@ A.2 Model Architecture
Building upon the Qwen3 (Yang et al., 2025) architectures, we perform our experiments utilizing both dense and MoE architectures. Notably, we adopt an auxiliary loss free approach (Liu et al., 2024) for the training of the MoE models. Detailed architecture settings are provided in Table 1, where $E$ denotes the total number of experts and $E_{a}$ denotes the number of active experts.
@@ A.3 Experiment Results
We report comprehensive evaluation results to demonstrate performance progression throughout the training process. Tables 2 to 11 present the pre-training results across various models and different training tokens. Similarly, Tables 12 to 17 summarize the performance metrics for the mid-training stage.
@@ B.1 Implementation Details
For RLVR on mathematical reasoning tasks, we employ the on-policy GRPO algorithm (Shao et al., 2024) without KL regularization. Following Yu et al. (2025), we incorporate clip-higher and dynamic sampling strategies to stabilize training. The process is conducted in two stages: an initial 700 steps with a sequence length of 8K, followed by continued training at a sequence length of 16K. We maintain a batch size of 128 and a constant learning rate of $1\times 10^{-6}$ for two stages. During training, we sample 16 outputs per prompt at a temperature of 1.0.